OpenStudy (anonymous):

gerald and michelle went on a 24 mile bike ride. they rode 1/4 of the distance before stopping. by lunchtime, they rode 5/8 of the total distance. how many miles did they have left to ride after lunch?

OpenStudy (anonymous):

1 - 1/4 - 5/8 = 1/8 24 x 1/8 = 3 miles left

OpenStudy (anonymous):

\[\frac{3}{8}24=9 \text{ miles} \]

OpenStudy (anonymous):

can u tell me how you got 3/8

OpenStudy (anonymous):

"by lunchtime, they rode 5/8 of the total distance." is the key. 8/8 is the whole distance of 24. They had traveled 5/8 at lunch time. That leaves 8/8 - 5/8 = 3/8 left to ride.

OpenStudy (anonymous):

thanks that was helpful

OpenStudy (anonymous):

Your welcome.